Crew 216 Operations Report 05Dec19

[title Operations Report– December 05th]

Crew 216 Operations Report 05 Dec-19
SOL: 11
Name of person filing report: Mike Lawson
Non-nominal systems: NA
Generator: run
Hours run: 13.5
From what time last night: 1728
To what time this morning: 0657
List any additional daytime hours when the generator was run: N/A
Solar— SOC % (Before generator is run at night): 77%
Diesel Reading – 33
Station Propane Reading – 51
Ethanol Free Gasoline: N/A
Water (loft tank) (gal): 47
Water Meter (units): 1465338
Water (static tank) (gal): 184
Static to Loft Pump used – yes
Water in Green Hab (gal): 244
Water in Science Dome (gal): 0
Toilet tank emptied: yes
Deimos rover used: no
Beginning charge:
Ending charge:
Currently charging:
Sojourner rover used: no
Beginning charge:
Ending charge:
Currently charging:
Spirit rover used: no
Hours: 116
Beginning charge: (Before EVA): 100
Ending charge: (On return from EVA, before recharging): 100
Currently charging:
Opportunity rover used: no
Hours: 68.7
Beginning charge: 100
Ending charge: 100
Currently charging: yes
Curiosity rover used: no
Hours: 121
Beginning charge: 100
Ending charge: (On return from EVA, before charging): 100%
Currently charging: yes
Notes on rovers: Opportunity and Deimos off-site for maintenance.
ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): no
Reason for use: N/A
Oil Added? no
# Hours the ATVs were used today: 0
Notes on ATVs:
HabCar used and why, where?
CrewCar used and why, where?
General notes and comments: Weather day, no EVAs. Crew meeting to plan station standown/handoff/end of sim.
Summary of internet: Erratic/unstable wifi connection. Director reports possible issue with MagWeb wireless gateway.
Summary of suits and radios: Suit 4 airflow is weak, possibly a battery charging issue.
Summary of Hab operations: Routine maintenance and housekeeping tasks. Hot water line vibration is extreme when pump engages. Recommend securing it to HAB wall or risk fitting failure.
Summary of GreenHab operations: routine watering.
Summary of Science Dome operations: Test ingots of plaster of paris subjected to destructive load testing.
Summary of RAM operations: RAM Rover power charging stations inop, electrical issues remain unresolved.
Summary of any observatory issues: NA
Summary of health and safety issues: NA
Questions, concerns and requests to Mission Support: 1 bag B, 1 bag NB trash in airlock.


Copyright © The Mars Society. All rights reserved. | Main Site